OpenStudy (anonymous):

What is the solution to the equation 2 ^(1 - 2x) = 3

OpenStudy (anonymous):

you can start by log (2^(1-2x)) = log 3

OpenStudy (anonymous):

then the power would get down... \[\LARGE (1-2x)\log (2)=\log (3)\] \[\LARGE (1-2x)=\frac{\log (3)}{\log (2) }\] ...